Will rate!!

Will rate!!

media%2F108%2F10827c28-69e4-422b-95cc-bc
media%2F823%2F823ffbf7-3166-48bb-b8b5-4d
Answer #1

a) The highest value of P(x) is 0.208 for when X=$300. Hence that is the most likely dollar amount that can be won.

ans: The most likely dollar amount is $300

b) When you spin Lose a turn or Bankrupt, the amount won is x=$0. The probability value P(x) for x=0 is 0.083.

ans: The probability to spin Lose a turn or Bankrupt is 0.083

c) the average dollar amount is the expected value of X. It is calculated as below

E(x)-XP() = 0 × P(0) 300 × P(300) 350 × P(350) + 400 × P(400) 450 × P(450) + 500 × P(500) + 550 × P(550) 600 × P(600) 700 × P(700) 800 × P(800) + 900 × P(900) + 5000 × P(5000) = 0 x 0.083 + 300 × 0.208 + 350 x 0.042 + 400 x 0.083 + 450 x 0.042 + 500 x 0.125 + 550 x 0.042 + 600 x 0.125 + 700 x 0042 + 800 x 0,083 + 900 x 0.083 + 5000 х 0042 668.75

The average dollar amount is $668.75

d) The probability that the contestant spins $0 in any given turn is P(x=0)=0.083

the probability that the contestant spin $0 each time is

(Probability that spin $0 first time)x(Probability that spin $0 second time)x(Probability that spin $0 third time) =

P(X=0)xP(X=0)xP(X=0) = 0.083*0.083*0.083 = 0.0006.

We can also solve this using a Binomial distribution as given below.

Let Y be the number of times the contestant spins $0 out of 3 spins. We can say that y has a Binomial distribution with parameters, number of trials n=3 and the success probability p=0.083.

The probability of spinning $0, Y=y number of times is given by (the Binomial probability is)

0.083(1 0.083)3- 3! 3-10.083(1 - 0.083)

The probability that the contestant spin $0, 3 out of 3 times is given by the probability of Y=3

-C ) 0.0833(1 0.083)-a P(Y = 3) 3! 3!(3-370 0833 (1-0.083)0 = 0.0833 0.0006

ans: The probability that the contestant spin $0, 3 out of 3 times is 0.0006

e) The probability that the contestant spins $0 at least one time out of 3 is same as 1- (probability that the contestant does not spin $0 each of 3 times)

The probability that the contestant does not spin $0 in any given turn is 1-P(x=0)=(1-0.083) = 0.917

the probability that the contestant spin does not $0 each of 3 times is

(Probability that does not spin $0 first time)x(Probability that does not spin $0 second time)x(Probability that does not spin $0 third time) =(1-P(X=0))x(1-P(X=0))x(1-P(X=0)) = 0.917*0.917*0.917 = 0.7703

The probability that the contestant spins $0 at least one time out of 3 is same as 1- (probability that the contestant does not spin $0 all 3 times) = 1-0.7703=0.2297

Alternatively, we can use Binomial distribution.

The probability that the contestant spin $0 at least one time out of 3 is same as the probability that Y is 1 or more

P(Y > 1)=1-P(y < 1) -10.083(1 0.083)o 3-0 3! 03-OT0.083(1 0.083) 1-(1-0.083)3 1- 0.7703 0.2297